In 1999, the base tuition at Concordia University for an undergraduate Quebec resident was $1668.30 and in 2013, it was $2224.20. If the Quebec CPI (2002 = 200) increased from 93.5 to 121.7 over this period, calculate the real increase in the price of tuition.

The answer is: 2.4%.

Please show the detailed step-by-step work leading to this answer. Thank you very much.
Answer #1

CPI increased from 93.5 to 121.7. This means the price level has increased by ((121.7/93.5) - 1)x100 = 30.16%.

to calculate the real value of tuition fee in 2013 we need to discount $2224.2 by 30.16%.

Real value of tuition fee in 2013 = 2224.2/(1+0.3016) = 1709.28.

To determine real increase we need to find ((1709.28/1668.3)-1)x100 = 2.4%